OpenStudy (anonymous):

The sales tax in New york city is 8.25%. How much tax is charged on a purchase of 5 juicers at $73 apiece? what is the total price? the amount of sales tax is $....

OpenStudy (anonymous):

that's easy

OpenStudy (anonymous):

(5 * 73) * (8.25/100) = 30.1125

OpenStudy (anonymous):

Proportion. 8.25/100 = x/(5(73))

OpenStudy (anonymous):

total price = (5 * 73)

OpenStudy (anonymous):

total price = (5 * 73) + 30.1125

OpenStudy (anonymous):

ok thnx